------------- Begin Forwarded Message ------------- X-Sender: cwolf@mail.hst.nasa.gov Mime-Version: 1.0 Date: Wed, 25 Mar 1998 06:49:55 -0400 To: george.albright@hq.nasa.gov, sadams@estif2.gsfc.nasa.gov, barnbaum@stsci.edu, preston.m.burch.1@gsfc.nasa.gov, jconwell@ems.jsc.nasa.gov, john.h.campbell.1@gsfc.nasa.gov, bcooke@sssp.hst.nasa.gov, carpenter@tma1.gsfc.nasa.gov, aferro@as.arizona.edu, jeffrey.m.hanley1@jsc.nasa.gov, hst-status@stsci.edu, chenning@pop500.gsfc.nasa.gov, gsfc.hst.general@newsfeed.gsfc.nasa.gov, j.keith.kalinowski.1@gsfc.nasa.gov, ken.ledbetter@hq.nasa.gov, j_lesko@swales.com, jlegg@stx.com, jleibee@hst.nasa.gov, dleckrone@hst.nasa.gov, ann.c.merwarth.1@gsfc.nasa.gov, jmccullo@ems.jsc.nasa.gov, gmorrow@hst.nasa.gov, mgw@resource-intl.com, mnolan@pop100.gsfc.nasa.gov, Jan.Owings@gsfc.nasa.gov, kpulkkinen@v2pop.hst.nasa.gov, gilbert.roth@hq.nasa.gov, rrasche@as.arizona.edu, arankin@pop400.gsfc.nasa.gov, steven.l.smith1@jsc.nasa.gov, stdesk@eso.org, rthompson@as.arizona.edu, edward.weiler@hq.nasa.gov, vweyers@pop400.gsfc.nasa.gov, howard.j.wood.1@gsfc.nasa.gov From: Charlie Wolf Subject: Daily Report #2087 03/25/98 HUBBLE SPACE TELESCOPE DAILY REPORT #2087 PERIOD COVERED: 0000Z (UTC) 03/24/98 - 0000Z (UTC) 03/25/98 Daily Status Report as of 084/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.5 Completed NIC/1/2/3 7819 (A New Stellar Generation in 30 Doradus) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the 30-Doradus nebula. There is ample evidence for current star formation within this nebula, which appears to have been triggered by the massive central cluster {R136}, in and just beyond the curved nebular filaments surrounding it. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2 7264 (High Redshift Reflection Nebulae) The Near Infrared Camera (Cameras One and Two) was used to examine the spatially extended K band light from high redshift radio galaxy TX0830+191 which cannot be dominated by stars. This object is a giant reflection nebulae rather than a "galaxy". The observations will allow us to: {I} characterize the morphology of the reflection nebula and the associated polarization field, {ii} infer the distribution of the dust (e.g wind vs. dust halo) and some of the properties of the dust, {iii} place constraints on the underlying stellar population of the host object, and {iv} investigate the ramifications of these objects for galaxy formation and evolution. The observation completed normally with no reported problems. 1.7 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star LHS1876. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Two Sets of STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. During the second iteration of this proposal, about 45 minutes before the end of this two-hour proposal, a guide star acquisition failed as detailed in 2.1 and HSTAR 6407. From that point, the take data flag remained down. Otherwise, the observations were completed as planned, and no further problems were reported. 1.9 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/1/2/3 7837 (The Changeover From H to He Rich Mass Loss in PN) The Near Infrared Camera (Cameras One,Two and Three) was used to make observations of the bright planetary nebula NGC-5315. This planetary nebula has helium rich stellar atmospheres and helium poor nebulae. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 6804 (Large Population Studies of Globular Clusters) The WF/PC-2 was used to make observations of the globular cluster NGC-2808 in order to obtaining very accurate photometric data on a complete sample of Post-Main Sequence stars over the entire area of the cluster for a sample of a "prototype" Galactic Globular Cluster. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 The acquisition scheduled for 084/005502Z failed when the search radius was exceeded. The re-acquisiiton at 084/0220Z also failed. HSTAR 6407 was written. The proposal described in 1.8 was affected, as were WF/PC-2 proposal 7712 and NIC proposal 7596, both of which will be described in tomorrow's report. Scheduled Reacquisitions : 5 Successful : 4 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The EDAC error counter was cleared at 083/1229Z per ROP SR-01. Using ROP DF-07, the six ephemeris tables were uplinked at 083/1023Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH ------------- End Forwarded Message -------------